OpenStudy (anonymous):

solve using law of cosines: a=14, c=18, B=48

OpenStudy (anonymous):

b^2 = a^2 + c^2 -2ac cos(48)

OpenStudy (lalaly):

how is the angle 48?

OpenStudy (anonymous):

therefore: b^2 = 14^2 + 18^2 - 2*14*18*cos(48)

OpenStudy (anonymous):

Why not Lalaly? He means degrees

OpenStudy (anonymous):

the angle is 48 degrees, my apologies

OpenStudy (lalaly):

ok

OpenStudy (anonymous):

doctorwho, we gave you the equation, solve it now

OpenStudy (anonymous):

i need the rest, i can do it, i just dont have the time, its my last problem and is do in, well... shortly. please finish it?

OpenStudy (anonymous):

come on guys, just a little farther please!

OpenStudy (anonymous):

13.52